OpenStudy (anonymous):

Find the value of 3 times x over the square root of x cubed. 3 over x. 3 times the square root of x. 3 over the square root of x. 3 times the square root of x, all over x.

OpenStudy (ttop0816):

Find the value of 3 times x over the square root of x cubed. =3x / √x^3 =3x / x√x =3x / x√x x √x / √x =3x(√x) / x√x(√x) =3x√x / x√x^2 =3x√x / x(x) =3x√x / x^2 =3√x / x, 3 times the square root of x, all over x!

OpenStudy (anonymous):

thank you so much!!

OpenStudy (ttop0816):

your welcome! do you mind helping me out too??

OpenStudy (anonymous):

what do you need help on?